OpenStudy (love_to_love_you):

Please help! Will medal and fan!! and testimony! What is the equation, in standard form, of a parabola that contains the following points? Show all work. (-2, 20), (0, -4), (4, -20)

OpenStudy (love_to_love_you):

@confluxepic

OpenStudy (confluxepic):

OK.

OpenStudy (love_to_love_you):

? @confluxepic

OpenStudy (love_to_love_you):

@cwrw238

OpenStudy (love_to_love_you):

@Zale101

OpenStudy (love_to_love_you):

@perl

OpenStudy (love_to_love_you):

@surjithayer

OpenStudy (anonymous):

y=ax^2+bx+c (0, -4) -4=a(0)^2+b(0)+c -4=c (-2, 20) 20=a(-2)^2+b(-2)-4 20=4a-2b-4 16=4a-2b ----------------- (1) (4, -20) -20=a(4)^2+b(4)-4 -20=16a+4b-4 -16=16a+4b ----------------- (2) Solve equation 1 & 2 for a and b, and then put values of a,b,c in y = ax^2 + bx + c

OpenStudy (love_to_love_you):

can you help me solve them?

OpenStudy (love_to_love_you):

1)a= 1/2b+4 b=2a-8 2)a=-1/4b-1 b= -4a-4

OpenStudy (love_to_love_you):

is that right?

OpenStudy (love_to_love_you):

I am really bad at this...

OpenStudy (anonymous):

\[(16 = 4a - 2b) + (-16 = 16 a + 4b) \] \[(2(16 = 4a - 2b) ) + ( -16 = 16 a + 4b)\] (32 + (-16) )= (8 a + 16a) + (-4b + 4b) 16 = 32a 1/2 = a (-2, 20) 20=1/2(-2)^2+b(-2)-4 20=2 -2b-4 16= -2 -2b 2b = -18 b = -9 y = 1/2(x^2) -9x - 4